OpenStudy (anonymous):

A circular pool has radius 7.1 meters. You want to build a cement walkway all the way around it and you have enough cement for 33 square meters. How wide can you make your walkway? Give your answer to the nearest centimeter.

OpenStudy (amistre64):

take the circomference of the pool and see what that measures

OpenStudy (amistre64):

or...or...... you have the area of 2 circles to find, one insid ethe other

OpenStudy (amistre64):

Area larger circle - Area smaller circle = 33

OpenStudy (amistre64):

pi(R^2) - pi(r^2) = 33 r = 7.1

OpenStudy (amistre64):

pi(R^2 - (7.1)^2) = 33 solve for R

OpenStudy (anonymous):

I got \[\sqrt{61}\]

OpenStudy (amistre64):

you got? prove it :)

OpenStudy (anonymous):

lol ok hold on

OpenStudy (anonymous):

YAYY

OpenStudy (amistre64):

yay!! was we right?

OpenStudy (anonymous):

YES! :)

OpenStudy (anonymous):

you are soo much more helpful than my teacher

OpenStudy (anonymous):

i learn more from you than i do form 2 hours of class

OpenStudy (amistre64):

we must be what they wrote about when they defined genius in the dictionary :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

I forget more in 2 hours than I ever dreamed of learning :)